# Interpretable and flexible non-intrusive reduced-order models using reproducing kernel Hilbert spaces

### v1

@@ Appendix A Quadratic systems with QM approximations
$\qquad\qquad+\tilde{\mathbf{H}}_{4}({\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}+{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}+{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}})\big{)}\in{}^{r\times r^{3}},$
$\hat{\mathbf{H}}_{4}$ $={\boldsymbol{\Sigma}}^{-1}\tilde{\mathbf{H}}_{4}({\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}})\in{}^{r\times r^{4}}.$
@@ Appendix B Stability for linear systems
where $\tilde{\mathbf{A}}=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}$ and $\tilde{\mathbf{H}}=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}$. We then have a stability estimate for the ROM solution
@@
[…]
@@
$\tilde{\mathbf{q}}(t)^{\mathsf{T}}\frac{\textrm{d}}{\textrm{d}t}\tilde{\mathbf{q}}(t)&=\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]\\ &=\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}_{sym}\mathbf{V}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]\\ &\leq\lambda\left\|\mathbf{V}\tilde{\mathbf{q}}(t)\right\|_{2}^{2}+\left\|\mathbf{A}\right\|_{2}\left\|\mathbf{W}\right\|_{2}\left\|\mathbf{V}\tilde{\mathbf{q}}(t)\right\|_{2}\|\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\|_{2}\\ &=\lambda\left\|\tilde{\mathbf{q}}(t)\right\|_{2}^{2}+\left\|\mathbf{A}\right\|_{2}\left\|\mathbf{W}\right\|_{2}\left\|\tilde{\mathbf{q}}(t)\right\|_{2}\|\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\|_{2},\end{aligned}$ (B.4)
@@
Appendix C Main nomenclature
@@
$K:{}^{n_{x}}\times{}^{n_{x}}\to\real$ symmetric kernel function
@@
$\mathbf{x}_{j}\in{}^{n_{x}}$ inputs for kernel interpolation
$\mathbf{y}_{j}\in{}^{n_{y}}$ outputs for kernel interpolation
$\mathbf{v}:{}^{n_{x}}\to{}^{n_{y}}$ function to interpolate: $\mathbf{y}_{j}=\mathbf{v}(\mathbf{x}_{j})$
@@
${\boldsymbol{\Omega}}\in{}^{m\times n_{y}}$ coefficient matrix for kernel interpolation
@@
${\boldsymbol{\psi}}_{\!\epsilon}:{}^{n_{x}}\to{}^{m}$ RBF kernel evaluation function
${\boldsymbol{\phi}}:{}^{n_{x}}\to{}^{n_{\phi}}$ feature map
$\mathbf{G}\in{}^{n_{\phi}\times n_{\phi}}$ weighting matrix for feature map kernels
$\mathbf{C}\in{}^{n_{y}\times n_{\phi}}$ post-feature map kernel coefficients
@@
$\mathbf{q}(t)\in{}^{n_{q}}$ full-order model state
$\mathbf{f}:{}^{n_{q}}\to{}^{n_{q}}$ full-order model dynamics function
${\boldsymbol{\mu}}\in{}^{n_{\mu}}$ parameters for the initial condition
$\mathbf{Q}\in{}^{n_{q}\times M(n_{t}+1)}$ shifted state snapshot matrix (all trajectories)
@@
$\tilde{\mathbf{q}}(t)\in{}^{r}$ intrusive reduced-order model state
$\hat{\mathbf{q}}(t)\in{}^{r}$ non-intrusive reduced-order model state
$\mathbf{V}\in{}^{n_{q}\times r}$ proper orthogonal decomposition (POD) basis matrix
$\mathbf{W}\in{}^{r\times r(r+1)/2}$ quadratic manifold (QM) weight matrix
$\mathbf{g}:{}^{r}\to{}^{n_{q}}$ decompression map
$\mathbf{h}:{}^{n_{q}}\to{}^{r}$ compression map

### v2

@@ Appendix A Quadratic systems with QM approximations
$\qquad\qquad+\tilde{\mathbf{H}}_{4}({\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}+{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}+{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}})\big)\in\real^{r\times r^{3}},$
$\hat{\mathbf{H}}_{4}$ $={\boldsymbol{\Sigma}}^{-1}\tilde{\mathbf{H}}_{4}({\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}})\in\real^{r\times r^{4}}.$
@@ Appendix B Stability for linear systems
where $\tilde{\mathbf{A}}=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}$ and $\tilde{\mathbf{H}}=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}$. We then have a stability estimate for the ROM solution.
@@
$\left\|\tilde{\mathbf{q}}(t)\right\|_{2}\leq\left\|\mathbf{A}\right\|_{2}\left\|\mathbf{W}\right\|_{2}\int_{0}^{t}\|\tilde{\mathbf{q}}(s)\otimes\tilde{\mathbf{q}}(s)\|_{2}e^{\lambda(t-s)}ds+e^{\lambda t}\left\|\mathbf{V}^{\mathsf{T}}\mathbf{q}_{0}\right\|_{2}.$ (B.3)
@@
$\begin{aligned} \tilde{\mathbf{q}}(t)^{\mathsf{T}}\frac{\textrm{d}}{\textrm{d}t}\tilde{\mathbf{q}}(t)&=\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]\\ &=\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}_{sym}\mathbf{V}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{q}}(t)^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]\\ &\leq\lambda\left\|\mathbf{V}\tilde{\mathbf{q}}(t)\right\|_{2}^{2}+\left\|\mathbf{A}\right\|_{2}\left\|\mathbf{W}\right\|_{2}\left\|\mathbf{V}\tilde{\mathbf{q}}(t)\right\|_{2}\|\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\|_{2}\\ &=\lambda\left\|\tilde{\mathbf{q}}(t)\right\|_{2}^{2}+\left\|\mathbf{A}\right\|_{2}\left\|\mathbf{W}\right\|_{2}\left\|\tilde{\mathbf{q}}(t)\right\|_{2}\|\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\|_{2},\end{aligned}$ (B.4)
@@
Appendix C Error estimate between intrusive ROM and Kernel ROM
We derive an error result comparing the intrusive projection-based ROM solution $\tilde{\mathbf{q}}(t)$ and the Kernel ROM solution $\hat{\mathbf{q}}(t)$. Let $\hat{\mathbf{e}}(t)=\tilde{\mathbf{q}}(t)-\hat{\mathbf{q}}(t)$, which satisfies the ODE
$\frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{e}}(t)=\tilde{\mathbf{f}}(\tilde{\mathbf{q}}(t))-\hat{\mathbf{f}}(\hat{\mathbf{q}}(t)),\qquad\hat{\mathbf{e}}(0)=\mathbf{0}.$ (C.1)
We then have the following.
Proposition C.1.
Let $\hat{\mathbf{M}}\in\real^{r\times r}$ be a symmetric positive definite weighting matrix with Cholesky factorization $\hat{\mathbf{M}}=\hat{\mathbf{L}}\hat{\mathbf{L}}^{\mathsf{T}}$. If $\hat{\mathbf{f}}$ is an unregularized kernel interpolant of $\tilde{\mathbf{f}}+{\boldsymbol{\delta}}\in{\cal H}_{K}^{r}$ where $\left\|{\boldsymbol{\delta}}(\hat{\mathbf{q}}(s))\right\|_{\hat{\mathbf{M}}}<\delta(s)$, then
$\left\|\hat{\mathbf{e}}(t)\right\|_{\hat{\mathbf{M}}}$ $\leq\int_{0}^{t}\left(P_{K,\tilde{\mathbf{Q}}}(\hat{\mathbf{q}}(s))\|\hat{\mathbf{L}}\|_{2}\|\tilde{\mathbf{f}}+{\boldsymbol{\delta}}\|_{{\cal H}_{K}^{r}}+\delta(s)\right)e^{\int_{s}^{\mathsf{T}}\Lambda_{\hat{\mathbf{M}}}[\hat{\mathbf{f}}](\hat{\mathbf{q}}(\tau))d\tau}ds,$ $\forall\;t$ $\in(0,T).$ (C.2)
Proof.
The dynamics in eq. C.1 can be rewritten as
Therefore
Applying Lemma 5.1 yields the result. ∎
Appendix D Main nomenclature
@@
$K:\real^{n_{x}}\times\real^{n_{x}}\to\real$ symmetric kernel function
@@
$\mathbf{x}_{j}\in\real^{n_{x}}$ inputs for kernel interpolation
$\mathbf{y}_{j}\in\real^{n_{y}}$ outputs for kernel interpolation
$\mathbf{v}:\real^{n_{x}}\to\real^{n_{y}}$ function to interpolate: $\mathbf{y}_{j}=\mathbf{v}(\mathbf{x}_{j})$
@@
${\boldsymbol{\Omega}}\in\real^{m\times n_{y}}$ coefficient matrix for kernel interpolation
@@
${\boldsymbol{\psi}}_{\!\epsilon}:\real^{n_{x}}\to\real^{m}$ RBF kernel evaluation function
${\boldsymbol{\phi}}:\real^{n_{x}}\to\real^{n_{\phi}}$ feature map
$\mathbf{G}\in\real^{n_{\phi}\times n_{\phi}}$ weighting matrix for feature map kernels
$\mathbf{C}\in\real^{n_{y}\times n_{\phi}}$ post-feature map kernel coefficients
@@
$\mathbf{q}(t)\in\real^{n_{q}}$ full-order model state
$\mathbf{f}:\real^{n_{q}}\to\real^{n_{q}}$ full-order model dynamics function
${\boldsymbol{\mu}}\in\real^{n_{\mu}}$ parameters for the initial condition
$\mathbf{Q}\in\real^{n_{q}\times M(n_{t}+1)}$ shifted state snapshot matrix (all trajectories)
@@
$\tilde{\mathbf{q}}(t)\in\real^{r}$ intrusive reduced-order model state
$\hat{\mathbf{q}}(t)\in\real^{r}$ non-intrusive reduced-order model state
$\mathbf{V}\in\real^{n_{q}\times r}$ proper orthogonal decomposition (POD) basis matrix
$\mathbf{W}\in\real^{r\times r(r+1)/2}$ quadratic manifold (QM) weight matrix
$\mathbf{g}:\real^{r}\to\real^{n_{q}}$ decompression map
$\mathbf{h}:\real^{n_{q}}\to\real^{r}$ compression map